# Can Language Models Reason about

### v1

@@ 3. Format Your Response: Please provide your response in the following format: [Your Response]:
B.2 Probing Results
Refined vs. Polar value-expressing statements. We experiment with using refined value-expressing statements (e.g., “I strongly agree…” vs. “I somewhat agree…”) instead of polar statements (e.g., “I agree…” vs. “I disagree…”) as demonstrations to LMs. Table 14 shows that refined statements prove more effective in aiding language models to make predictions, underscoring the importance of precise and nuanced value expressions.
Probing results broken down by three probe setups. Table 15 shows the results of the probing experiments under the polar evaluation scheme broken down by the three probing sets, corresponding to the main probe results in Figure 3.
Breakdown $\sigma$Inequity scores of all probed models. Full results of $\sigma$Inequity of all probed models per each of the considered demographics dimensions are shown in Table 16.
How do different types of statement influence the prediction of the other types? Figure 8 illustrates how using different categories of value statements as demonstrations affects the prediction of other categories. Our results indicate that value statements are not limited to strongly predicting only within their own category; in some cases, other categories can perform surprisingly well in predicting different types of value choices. This finding highlights intriguing dynamics and connections between various categories of value statements.
The uneven individualistic value reasoning ability of GPT-4o across demographics groups. Figure 9 shows the performance disparity across demographic groups of different demographic dimensions.
How do demographic statements impact weak models like GPT-4o-mini in individualistic value reasoning? Figure 10 compares probing setups with and without demographic information with GPT-4o-mini. For such a weaker model, including demographics leads to significantly better predictions than providing value statements alone, as the model is likely to struggle to interpret nuanced descriptive value statements compared to direct demographic identity deceleration.
Demonstration Probe 0 Probe 1 Probe 2 Average
@@ B.2 Probing Results
Appendix C Details of the Individualistic Value Reasoner
C.1 Training Setups
@@
Option 1: I believe that claiming government benefits
to which you are not entitled is not justifiable

### v2

@@ 3. Format Your Response: Please provide your response in the following format: [Your Response]:
C.2 Probing Results
Refined vs. Polar value-expressing statements.
We experiment with using refined value-expressing statements (e.g., “I strongly agree…” vs. “I somewhat agree…”) instead of polar statements (e.g., “I agree…” vs. “I disagree…”) as demonstrations to LMs. Table 13 shows that refined statements prove more effective in aiding language models to make predictions, underscoring the importance of precise and nuanced value expressions.
Probing results broken down by three probe setups.
Table 14 shows the results of the probing experiments under the polar evaluation scheme broken down by the three probing sets, corresponding to the main probe results in Figure 2.
Breakdown $\sigma$Inequity scores of all probed models.
Full results of $\sigma$Inequity of all probed models per each of the considered demographics dimensions are shown in Table 15.
How do different types of statement influence the prediction of the other types?
Figure 8 illustrates how using different categories of value statements as demonstrations affects the prediction of other categories. Our results indicate that value statements are not limited to strongly predicting only within their own category; sometimes, other categories can perform surprisingly well in predicting different types of value choices. This finding highlights intriguing dynamics and connections between various categories of value statements.
The uneven individualistic value reasoning ability of GPT-4o across demographics groups.
Figure 9 shows the performance disparity across demographic groups of different demographic dimensions.
How do demographic statements impact weak models like GPT-4o-mini in individualistic value reasoning?
Figure 10 compares probing setups with and without demographic information with GPT-4o-mini. For such a weaker model, including demographics leads to significantly better predictions than providing value statements alone, as the model is likely to struggle to interpret nuanced descriptive value statements compared to direct demographic identity deceleration.
Demo. Probe 0 Probe 1 Probe 2 Average
@@ C.2 Probing Results
Appendix D Details of the Individualistic Value Reasoner
D.1 Training Setups
@@
    Option 1: I believe that claiming government benefits to
    which you are not entitled is not justifiable
@@ D.2 Additional Results
Polar Refined All
Method Probe 1 Probe 2 Probe 3 Avg. Probe 1 Probe 2 Probe 3 Avg. Avg.
Random 46.37 45.51 44.23 45.37 29.16 29.03 25.43 27.87 36.62
Resemble (1600, top 1) 66.41 65.95 64.59 65.65 47.03 48.07 47.55 47.55 56.60
Resemble (1600, top cluster) 71.79 71.36 68.01 70.39 54.02 55.79 53.27 54.36 62.38
Resemble (all, top 1) 70.31 70.15 69.02 69.83 53.26 54.01 53.27 53.51 61.67
Resemble (all, top cluster) 74.74 74.87 71.60 73.73 59.32 60.78 58.32 59.47 66.60
@@
Appendix E Acknowledgement of AI Assistance
We use AI solely for assistance with the language of the paper.